OpenStudy (anonymous):

Find the area of a regular hexagon with the given measurement. 48-inch perimeter A = sq. in.

OpenStudy (sriram):

side=48/6=8 area=8*8*(sqrt3)/4*6 166.27

OpenStudy (anonymous):

166

OpenStudy (anonymous):

166